# CoinFT: A Coin-Sized, Capacitive 6-Axis Force Torque Sensor for Robotic Applications

### v2

@@ III-D2 Sensitivity and Dynamic Range Tuning
The effect of varying pillar diameter on sensor response is illustrated in Fig. 4 (d)-(g). Both experimental and FEA results demonstrate a gradual decrease in sensitivity under normal load (Fig. 4 (d), (g)). This reduction in sensitivity is attributed to the stiffening of the pillars as they compress. The lateral deformation of the pillars (Fig. 4 (h)) increase their effective Young’s modulus, as described in equation (2). Given that the sensitivity of the capacitance change to further compression increases as shown in equation (1) and demonstrated in (Fig. 4 (g)), we note that the stiffening effect is significant. For instance, the quadratic increase of stiffness and sensitivity of capacitance offset each other at lower displacements, resulting in an approximately linear capacitance–force relationship (Fig. 4 (d), (g)). However, the stiffening effect outweighs the increase in capacitance change with higher deformation, explaining the decrease in sensitivity observed in Fig. 4 (d).
@@ III-D4 Robustness
Many robot applications involve incidental large impulses that can potentially damage commercially available 6-axis F / T sensors. Due to the compliance of the pillar arrays, CoinFT demonstrates reliable robustness against impacts. As shown in Fig. 7, CoinFT can accurately measure small loads of 1 N, 2 N, and 3 N even after sustaining an impact from a hammer that causes a force reading of 180 N. Such an impact would typically damage commercial sensors like the Gamma (ATI Industrial Automation) used for the CoinFT calibration. This level of robustness enables a broader range of robotic interactions beyond controlled environments, making the integration of 6-axis F / T sensors practical for real-world applications.
Figure 7: CoinFT is robust. It can produce reliable force readings after being struck by a hammer.
IV Contact Force Modulation of Aerial Vehicles
Drones offer unique capabilities for performing contact-rich tasks in inaccessible areas, such as collecting plant samples from cliffs in Hawaii [6] or cleaning windows on tall buildings [46]. However, most drones lack contact sensing due to the fragility and cost of existing load cells. Integrating robust and affordable force sensing would greatly enhance drone utility. CoinFT addresses this challenge with its compact design, lightweight construction (6.7 g including the sensor and USB-to-serial bridge), and affordability (under \$10 per unit). CoinFT is also durable enough to withstand impacts common in real-world applications. In this section, we present a drone equipped with CoinFT and a force controller, enabling the drone to attach a package of electronics to a horizontal surface. This force-controlled interaction, unlocked by CoinFT, allows drones to perform tasks such as deploying sensor packages for environmental surveillance or conducting structural health monitoring on tall buildings. These sensor packages can continue to operate well beyond the limited flight time of the drones, providing valuable data to humans.
IV-A Drone with CoinFT
@@
IV-B Attitude Controller for Contact
@@
Given that the focus here is control with respect to contact, we simplify the navigation aspect of the control problem by providing the drone with accurate estimates of its pose and orientation via an external motion capture system. We then use this as feedback to track a minimum snap trajectory spline [47] to a region near the contact surface. Crucially though, we do not provide the drone with the actual height of the contact surface. Instead, our policy will search through a range of heights within which we are confident that the contact surface exists and transform it into a sequence of desired positions ($\boldsymbol{p_{des}}$), velocities ($\boldsymbol{v_{des}}$) and orientations ($\boldsymbol{q_{des}}$), for our controller to track while it searches for contact.
@@
$\boldsymbol{K_{p}},\;\boldsymbol{K_{v}}=\begin{cases}\boldsymbol{K_{p}^{oc}},\;\boldsymbol{K_{v}^{oc}}&\text{if out of contact}\\ \boldsymbol{K_{p}^{ic}},\;\boldsymbol{K_{v}^{ic}}&\text{if in contact}\\ \end{cases}$
@@
$\boldsymbol{z_{cmd}}$ $=\frac{\boldsymbol{F_{des}}}{||\boldsymbol{F_{des}}||}$
$\boldsymbol{y_{cmd}}$ $=\frac{\boldsymbol{z_{cmd}\times}\boldsymbol{x_{des}}}{||\boldsymbol{z_{cmd}\times}\boldsymbol{x_{des}}||}$
@@
IV-C Force Controlled Deployment of Objects on Environment Surfaces
To test the designed force controller and perform electronics package delivery tasks, we conducted flight tests. The test setup includes a laser-cut acrylic plate horizontally attached to a structure made of extruded aluminum frames (Fig. 10 (a)), simulating an environmental surface. The integrated system consists of a ground control station (PC), motion capture system (OptiTrack, NaturalPoint, Inc.), and a drone equipped with a companion computer, flight controller, IMU, and a CoinFT sensor (Fig. 9). The proposed attitude controller runs on the companion computer, providing control inputs to the flight controller, which then converts these inputs into PWM signals for the four motors.
Figure 10: Results of the force controller test. (a) The drone ascends until contact is made with the test setup, where it initiates force control. (b) CoinFT readings and commanded normal force.
The performance of the proposed attitude-based force controller is illustrated in Fig. 10(b). A hemispherical contact tip (Fig. 8 (a)) is mounted on the drone. During the SEARCH state, the motor thrust ramps linearly until the drone senses contact. Contact is signaled by normal force reaching a desired value, which is initially set by the ground control station. Upon contact, the controller transitions to the HOLD state, where it tracks a pre-programmed sinusoidal force command. The force controller demonstrates accurate tracking, with an RMS error of 0.1779 N.
@@ V Discussion & Future Work
While this paper focuses primarily on demonstrating CoinFT’s capabilities on drone platforms, the sensor’s versatility extends to a wide range of other robotic applications. CoinFT has already been used in wearable haptic devices, where it measures and controls the contact force between the device and the human subject. Sarac et al. [48] used CoinFT to reveal that people perceive different intensities when the same force is applied in normal versus shear directions. Yoshida et al. [7] employed CoinFT to ensure consistency in user studies by initializing the haptic device tactor based on contact force rather than position. This approach is crucial for accommodating variations in users’ body shapes, as position-based initialization does not consistently deliver uniform haptic feedback. CoinFT’s slim and compact design also makes it highly suitable for wearable fingertip devices [43].
@@
While the current form factor of CoinFT has been designed to showcase its versatility across various robotic platforms such as drones, robot fingertips, and haptic devices, it can be tailored to meet the specific needs of other applications. For example, CoinFT can be made even narrower to sensorize compact medical devices [51]. The tradeoff in this case would be a decrease in sensitivity due to the reduced sensing area. Conversely, CoinFT could be scaled up to fit the form factor of quadruped or humanoid feet, allowing for the measurement of ground reaction forces (GRF). The larger sensing area in this configuration would enable both higher sensitivity and an expanded dynamic range. This would require a careful selection of design parameters to meet the specifications on sensitivity and dynamic range, as discussed in Section III-D.
While CoinFT is versatile and suitable for a wide range of robotic applications, it is not intended as a direct replacement for existing commercial sensors like the Gamma (ATI Industrial Automation). The mechanical bandwidth of CoinFT is an order of magnitude lower than that of the Gamma (1 kHz), making it less capable of detecting highly dynamic contacts. Additionally, the Gamma offers a finer force resolution even with its high dynamic response capabilities. Although CoinFT is highly robust against large compressive impulses, it is less resilient to large tensile loads, as the bonding between the silicone rubber pillars and the PCB may fail under such conditions.
Other challenges must be considered when using CoinFT. While shielding has been implemented to prevent the coupling of stray capacitance, the surface-mounted electrodes of the chip remain partially exposed, which can result in imperfect shielding. Furthermore, the silicone rubber pillars may creep under long-term sustained loads, altering the mechanical characteristics of the pillar array and necessitating re-calibration. For applications that require it, one could assemble CoinFT with a low-creep silicone elastomer [52] of similar stiffness.
The current thin PCB might not be entirely rigid and could deform out of plane under significant local pressure, potentially reducing calibration accuracy. This issue, however, can be mitigated by using a slightly thicker PCB or bonding a thin layer of stiff material. Additionally, variations in ambient temperature may cause slight deformations in the sensing electrodes or the internal electronics of the chip, leading to drift in the capacitance signal. Despite these limitations, CoinFT remains valuable for many robotic applications where such high precision and bandwidth in a 6-axis F / T sensor is not essential, and where robustness and low cost are prioritized.
@@ VI Conclusion
We demonstrated the utility of CoinFT by integrating it with a drone to perform force-controlled tasks, such as attaching an electronics package to environmental surfaces. This capability is an example of CoinFT’s potential to democratize the use of 6-axis F / T sensors across a broad range of robotic platforms, enabling more widespread adoption in applications that demand robustness and cost-effectiveness without sacrificing essential functionality. To promote this use of CoinFT technology, we have published solid models, fabrication instructions and software as an open-source design at https://hojung-choi.github.io/coinft.github.io.

### v3

@@ III-D2 Sensitivity and Dynamic Range Tuning
The effect of varying pillar diameter on sensor response is illustrated in Fig. 4 (d)-(g). Both experimental and FEA results demonstrate a gradual decrease in sensitivity under normal load (Fig. 4 (d), (g)). This reduction in sensitivity is attributed to the stiffening of the pillars as they compress. The constrained lateral deformation of the pillars (Fig. 4 (h)) increases their stiffness, as described in equation (2). Given that the sensitivity of the capacitance change to further compression increases as shown in equation (1) and demonstrated in (Fig. 4 (g)), we note that the stiffening effect is significant. For instance, the quadratic increase of stiffness and sensitivity of capacitance offset each other at lower displacements, resulting in an approximately linear capacitance–force relationship (Fig. 4 (d), (g)). However, the stiffening effect outweighs the increase in capacitance change with higher deformation, explaining the decrease in sensitivity observed in Fig. 4 (d).
@@ III-D4 Robustness
Many robot applications involve incidental large impulses that can potentially damage commercially available 6-axis F / T sensors. Due to the compliance of the pillar arrays, CoinFT demonstrates reliable robustness against impacts. As shown in Fig. 7 (a), CoinFT can accurately measure small loads of 1 N, 2 N, and 3 N even after sustaining an impact from a hammer that causes a force reading of 180 N. Such an impact would typically damage commercial sensors like the Gamma (ATI Industrial Automation) used for the CoinFT calibration. This level of robustness enables a broader range of robotic interactions beyond controlled environments, making the integration of 6-axis F / T sensors practical for real-world applications.
To further quantify the sensor’s impact resistance, a controlled drop test was conducted using a 0.53 kg chrome steel ball released from a height of 0.63 m, corresponding to an impact velocity of approximately 3.5 m / s and an impact energy of 3.3 J (Fig. 7 (b)). Before and after the ten repeated impacts, the CoinFT calibration was validated under a sinusoidal loading profile of 0–5 N at 0.5 Hz using a dual-mode muscle lever (Aurora Scientific Inc., 309C). The calibration performance remained consistent, exhibiting a mean-squared error of 0.055 N and an $R^{2}$ value of 0.982 when compared to the reference force. These results confirm that CoinFT maintains its measurement accuracy and structural integrity under repeated high-energy impacts.
While CoinFT exhibits strong robustness under compressive loading, it is less robust in the tensile direction, as excessive tensile stress can cause delamination between the silicone pillars and the PCB substrate. Tensile tests conducted on five CoinFT samples showed an average ultimate tensile strength of 2.84 MPa (143 N) with a standard deviation of 1.21 MPa (61 N). Increasing the pillar diameter can increase the absolute tensile force the sensor can withstand, while improved bonding strategies, such as surface treatments or adhesive modification, are expected to further enhance the ultimate tensile strength.
Figure 7: CoinFT robustness evaluation. (a) CoinFT continues to provide reliable force readings even after an impact from a hammer. (b) The calibration stays consistent after ten impacts from a 0.53 kg mass dropped at a height of 0.63 m.
Figure 8: Effects of temperature variation and its compensation. (a) Drift of raw signals due to temperature change. (b) Temperature compensation using an on-chip temperature sensor.
III-D5 Thermal Effects
We investigated the effect of temperature variation and implemented a compensation method using the on-chip temperature sensor of the PSoC. The thermal drift of the raw capacitance signals from all 12 channels was characterized in a temperature-controlled chamber. Across channels, the baseline capacitance drift was approximately 1–3% per 10,C. This drift profile was modeled using a second-order polynomial fit, achieving $R^{2}>0.9992$ for all channels. During inference, temperature compensation was performed by subtracting the predicted baseline drift from the raw capacitance signals prior to force/torque estimation. The ablation results under a no-load condition are shown in Fig. 8. With temperature compensation, the residual force error remained below 0.4 N and torque error below 1.5 mN·m, whereas without compensation, a clear temperature-dependent drift was observed in the force readings.
IV CoinFT Demonstrations
While numerous six-axis force/torque sensors have been developed, as discussed in Section II, few achieve a balance of compactness, slim form factor, low cost, accuracy, and robustness. This combination of features enables the broader adoption of F/T sensing across diverse robotic platforms. We demonstrate the versatility of CoinFT through a multi-axial contact probing setup that emulates robot fingertip interactions and a example that motivates its integration on drones.
IV-A Multi-axial Sensing for Robot Manipulators
Dexterous manipulation tasks, such as object handling or tool use, often require accurate multi-axial force/torque feedback, particularly when interactions require significant contact forces for success[57]. For integration on robot fingers, the sensing module must also maintain a compact and slim form factor to avoid drastically altering the hand geometry.
To evaluate CoinFT’s performance in such settings (e.g. as in [36]), we mounted a CoinFT beneath a robot fingertip with a hemispherical dome and applied varying force and torque along multiple axes using a 7-DoF robotic arm (Flexiv Rizon 4). As shown in Fig. 9 (a), a reference force/torque sensor (ATI Gamma) was positioned beneath CoinFT to provide ground-truth measurements. The resulting data (Fig. 9 (b)) demonstrate that CoinFT achieves accuracies listed in Table IV.
TABLE IV: Multi-axial sensing during contact probing.
Fx Fy Fz Mx My Mz
RMSE 0.108 0.120 0.378 5.669 3.874 0.720
$R^{2}$ 0.994 0.995 0.999 0.993 0.996 0.986
[…]
Figure 9: Multi-axial contact probing experiment. (a) The 7DoF arm applies force and torque in different axes on a hemispherical fingertip equipped with CoinFT, approximating manipulation scenarios. (b) CoinFT and reference sensor reading comparisons. Refer to Table IV for quantitative evaluation.
IV-B Contact Force Modulation of Aerial Vehicles
Drones offer unique capabilities for performing contact-rich tasks in environments that are difficult or dangerous for humans, such as collecting biological samples from cliffs [6] or cleaning vertical surfaces [58]. However, most drones lack reliable contact sensing due to the fragility, weight, and high cost of conventional force sensors. Without such feedback, maintaining stable contact or regulating interaction forces is challenging, often resulting in unsafe or inconsistent behaviors.
CoinFT directly addresses these limitations through its compact (20 mm diameter, 2 mm thickness), lightweight (2 g, or 6.7 g including wiring and electronics), and impact-tolerant design. In this section, we demonstrate that a drone equipped with CoinFT and a simple PID-based force controller can safely and stably perform contact interactions, such as pressing and attaching a payload onto a surface. This experiment highlights the ease of integrating CoinFT on aerial platforms; given its low cost and mechanical resilience, the sensor can withstand or be inexpensively replaced after crashes that are common in drone operations.
While only normal force feedback was utilized in this demonstration, CoinFT’s six-axis sensing capability may enable a broader range of aerial manipulation tasks, such as sliding, scraping, or compliant alignment that would not be feasible without multi-axial sensing.
IV-B1 Drone with CoinFT
@@
IV-B2 Attitude Controller for Contact
@@
Given that the focus here is control with respect to contact, we simplify the navigation aspect of the control problem by providing the drone with accurate estimates of its pose and orientation via an external motion capture system. We then use this as feedback to track a minimum snap trajectory spline [59] to a region near the contact surface. We do not provide the drone with the actual height of the contact surface. Instead, our policy will search through a range of heights within which we are confident that the contact surface exists and transform it into a sequence of desired positions ($\boldsymbol{p_{des}}$), velocities ($\boldsymbol{v_{des}}$) and orientations ($\boldsymbol{q_{des}}$), for our controller to track while it searches for contact.
@@
[…]
@@
[…]
@@
IV-B3 Force Controlled Deployment of Objects on Environment Surfaces
To test the designed force controller and perform electronics package delivery tasks, we conducted flight tests. The test setup includes a laser-cut acrylic plate horizontally attached to a structure made of extruded aluminum frames (Fig. 12 (a)), simulating an environmental surface. The integrated system consists of a ground control station (PC), motion capture system (OptiTrack, NaturalPoint, Inc.), and a drone equipped with a companion computer, flight controller, IMU, and a CoinFT (Fig. 11). The proposed attitude controller runs on the companion computer at 20 Hz, providing control inputs to the flight controller, which then converts these inputs into PWM signals for the four motors.
Figure 12: Results of the force controller test. (a) The drone ascends until contact is made with the test setup, where it initiates force control. (b) CoinFT readings and commanded force.
The performance of the proposed attitude-based force controller is illustrated in Fig. 12(b). A hemispherical contact tip (Fig. 10 (a)) is mounted on the drone. During the FREE state, the drone ascends until contact is sensed. In the SEARCH state, the thrust is linearly ramped up until a desired contact force is reached, which is initially set by the ground control station. The controller then transitions to the HOLD state, where it tracks a pre-programmed sinusoidal force command. The force controller demonstrates accurate tracking, with an RMS error of 0.1779 N. During flight tests, CoinFT showed no noticeable increase in noise or degradation in signal quality despite the potential electromagnetic interference from the motors [60], thanks to the effectiveness of its shielding design.
@@ V Discussion & Future Work
While this paper demonstrates CoinFT’s capabilities in contact probing or drone platforms, the sensor’s versatility extends to a wide range of other robotic applications. CoinFT has already been used in wearable haptic devices, where it measures and controls the contact force between the device and the human subject. Sarac et al. [35] used CoinFT to reveal that people perceive different intensities when the same force is applied in normal versus shear directions. Yoshida et al. [7] employed CoinFT to ensure consistency in user studies by initializing the haptic device tactor based on contact force rather than position. This approach is crucial for accommodating variations in users’ body shapes, as position-based initialization does not consistently deliver uniform haptic feedback. CoinFT’s slim and compact design also makes it highly suitable for wearable fingertip devices [54].
@@
While the current form factor of CoinFT has been designed to showcase its versatility across various robotic platforms such as drones, robot fingertips, and haptic devices, it can be tailored to meet the specific needs of other applications. For example, CoinFT can be made even narrower to sensorize compact medical devices [62]. The tradeoff in this case would be a decrease in sensitivity due to the reduced sensing area. Conversely, CoinFT could be scaled up to fit the form factor of quadruped or humanoid feet, allowing for the measurement of ground reaction forces (GRF). The larger sensing area in this configuration would enable both higher sensitivity and an expanded dynamic range. This would require a careful selection of design parameters to meet the specifications on sensitivity and dynamic range, as discussed in Section III-D. Moreover, the dual-mode electrode-switching strategy used in CoinFT may also inspire future capacitive tactile sensors for enhancing multi-axial sensitivity, particularly those employing programmable capacitance sensing microcontrollers that permit low-level reconfiguration of electrode connections.
While CoinFT is versatile and suitable for a wide range of robotic applications, it is not intended as a direct replacement for existing commercial sensors like the Gamma (ATI Industrial Automation). The mechanical bandwidth of CoinFT is an order of magnitude lower than that of the Gamma (1 kHz), making it less capable of detecting highly dynamic contacts. Additionally, the Gamma offers a finer force resolution even with its high dynamic response capabilities.
Other challenges must be considered when using CoinFT. While shielding has been implemented to prevent the coupling of stray capacitance, the surface-mounted electrodes of the chip remain partially exposed, which can result in imperfect shielding. While elastomer creep can affect signal, its effect is expected to be small in most applications. The silicone rubber used in CoinFT is a tin-cured RTV elastomer with mechanical properties similar to that in [48]. As noted then, it has low hysteresis and creep in comparison to other elastomers (e.g. urethanes) of similar stiffness. More detailed characterization of similar silicone materials can be found in [63, 49, 50, 64] For example, the material in [64] showed reversible viscoelastic recovery under loads up to 0.5 MPa, which would correspond to $\sim$32 N on CoinFT—suggesting that no permanent creep or damage occurs after repeated loading. Consequently, in applications that do not sustain a large load for long periods of time, mechanical creep is expected to contribute negligibly to signal drift, and recalibration or sensor replacement will not be required.
The current thin PCB might not be entirely rigid and could deform out of plane under significant local pressure, potentially reducing calibration accuracy. This issue, however, can be mitigated by using a slightly thicker PCB or bonding a thin layer of stiff material. Despite the limitations, CoinFT remains valuable for many robotic applications where such high precision and bandwidth in a 6-axis F / T sensor is not essential, and where robustness and low cost are prioritized.
@@ VI Conclusion
We demonstrated the utility of CoinFT by a multi-axial contact probing experiment emulating figertip-scale interactions, and integrating it with a drone to perform force-controlled tasks, such as attaching an electronics package to environmental surfaces. This capability is an example of CoinFT’s potential to democratize the use of 6-axis F / T sensors across a broad range of robotic platforms, enabling more widespread adoption in applications that demand robustness and cost-effectiveness without sacrificing essential functionality. To promote this use of CoinFT technology, we have published solid models, fabrication instructions and software as an open-source design at https://coin-ft.github.io/.